Please help me solve this: 5(2t-2)=30
Answer by elizabeth2005(20)

you start out by multiplying 5 by everything inside the par.
so 5 times 2x will be 10x
and 5 times 2 is 10
so now you have...10t-10=30

now you will add ten to both sides 10t-10=30
+10+10
---------
that then leaves you with 10t=40
then you divide both sides by 10 ten divided by ten cancels out and leaves you alone with the variable t...
ten divided by 40 is 4
so t= 4
after you find out what t is you should check your work
multiply 10 by four and subtract 10 that will leave you with thirty